In a random sample of 500 observations, we found the proportion of successes to be 48%.

Estimate with 95% confidence the population proportion of successes.


a) Repeat part (a) with n=200.

b) Repeat part (a) with n = 1000.

c) Describe the effect on the confidence interval estimate of increasing the sample size.


1
Expert's answer
2021-03-16T08:19:03-0400

Solution:

Given, "p=48\%=0.48, n=500"

So, "q=1-0.48=0.52"

(a) Standard error, "SE=\sqrt{\dfrac{pq}{n}}=\sqrt{\dfrac{0.48\times0.52}{500}}=0.022342"

z-value for 95% confidence interval is 1.96.

Now, confidence interval: "1.96\pm0.022342"

"=[1.937658,\ 1.982342]"

(b) n = 200

Standard error, "SE=\sqrt{\dfrac{pq}{n}}=\sqrt{\dfrac{0.48\times0.52}{200}}=0.035327"

z-value for 95% confidence interval is 1.96.

Now, confidence interval: "1.96\pm0.035327"

"=[1.924673,\ 1.995327]"

(c) n = 1000

Standard error, "SE=\sqrt{\dfrac{pq}{n}}=\sqrt{\dfrac{0.48\times0.52}{1000}}=0.015798"

z-value for 95% confidence interval is 1.96.

Now, confidence interval: "1.96\pm0.015798"

"=[1.944202,\ 1.975798]"

(d) Significant effect of increasing sample size on confidence interval is that it becomes precise or shorter than earlier.